# Neural network learns low-dimensional polynomials with SGD near the information-theoretic limit

### v1

@@ B.6 Second Layer Training
$!\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}$ $\lesssim(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda\delta_{0}}\mathopen{}\mathclose{{}\left(2N^{2}Q(R_{w},b,q^{\prime})+\mathbb{E}_{\boldsymbol{x}}[(f_{*})^{4}]}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2},$ (B.148)
with probability $1-4\delta_{0}$. Thus, since $\|\boldsymbol{a}^{*}\|^{2}=\tilde{O}(N)$, by setting $T_{2}=\tilde{\Theta}((N^{4}Q(R_{w},b,q^{\prime})+\mathbb{E}[f_{*}(\boldsymbol{x})^{4}])\varepsilon^{-2})$, and $N=\tilde{\Theta}(\varepsilon^{-1})$, we obtain that (B.136)$\lesssim\varepsilon$. ∎
Finally, we provide the approximation guarantee: If $\sigma$ is a degree-$q$ polynomial, we have the following result, which follows Lemmas 29 and 30 of [OSSW24].
@@
$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma\big{(}{\boldsymbol{w}_{j}^{2T_{1}}}^{\top}\boldsymbol{x}+b_{j}\big{)}-h(\boldsymbol{\theta}^{\top}\boldsymbol{x})}\right|=\tilde{O}(N^{-1}+\varepsilon)$ (B.149)
@@
We rely on the following result.
@@
Proof.  When $g_{q}(s)=\sigma(s)$ is a degree-$q$ polynomial,
$g_{q}(s)=\int_{b=-q}^{0}\sigma(s+b)\mathrm{d}b$ (B.151)
is also a degree-$q$ polynomial. Let us repeatedly define
$g_{q-i}(s):=g_{q-(i-1)}(s+1)-g_{q-(i-1)}(s)\quad(i=1,2,\cdots,q),$ (B.152)
and let $(c_{i,j})$ be coefficients so that $(s-1)^{i}=\sum_{j=0}^{i}c_{i,j}s^{j}$ holds for all $z$. Then, by induction, $g_{i}(s)$ is a degree-$i$ polynomial. Moreover, we have
$g_{q-i}(s)$ $=\sum_{j=0}^{i}c_{i,j}\int_{b=-q}^{0}\sigma(s+b+j)\mathrm{d}b$ (B.153)
$=2C_{b}\mathbb{E}_{b\sim\mathrm{Unif}([-C_{b},C_{b}])}\bigg{[}\bigg{(}\sum_{j=0}^{i}c_{i,j}\mathbbm{1}[j-q\leq b\leq j]\bigg{)}\sigma(s+b)\bigg{]},$ (B.154)
when $C_{b}\geq q$. Therefore, for any polynomial $h(s)$ with its degree at most $q$, there exists $\bar{v}(b;h)$ with $|\bar{v}(b;h)|\lesssim C_{b}$ such that for all $s$,
$\mathbb{E}[\bar{v}(b;h)\sigma(\delta s+b)]=h(s).$ (B.155)
Proof of Lemma 15. We now discretize Lemma 16. We focus on $N^{\prime}$ neurons that satisfy $\|\boldsymbol{w}_{j}^{2T_{1}}-\boldsymbol{\theta}\|\leq\varepsilon$ (by letting $a_{j}=0$ otherwise).
For $A=\tilde{\Theta}(N^{\prime})=\tilde{\Theta}(N)$ (with a small hidden constant), we consider $2A$ intervals $[-C_{b},C_{b}(-1+\frac{1}{A})),[C_{b}(-1+\frac{1}{A}),C_{b}(-1+\frac{2}{A})),\cdots,[C_{b}(1-\frac{1}{A}),C_{b}]$. By taking the hidden constant sufficiently small, for each interval there exists at least one $b_{j}$. Then, for $b_{j}$ corresponding to $[C_{b}(-1+\frac{i}{A}),C_{b}(-1+\frac{i+1}{A}))$, we set $a_{j}=\frac{N}{2}\int_{C_{b}(-1+\frac{i}{A})}^{C_{b}(-1+\frac{i+1}{A}))}\bar{v}(b;h)\mathrm{d}b$. Here we note that $|a_{j}|=\tilde{O}(1)$ holds for all $j$. If each interval contains more than one $b_{j}$, we ignore all but one by letting $a_{j}=0$ except for the one. By doing so, because of Lipschitzness of $\sigma$, we have
$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma(s+b_{j})-h(s)}\right|=\tilde{O}(N)$ (B.156)
for all $s=\tilde{O}(1)$. Because $|\boldsymbol{\theta}^{\top}\boldsymbol{x}^{t}|=\tilde{O}(1)$ with high probability, we have
$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma(\boldsymbol{\theta}^{\top}\boldsymbol{x}+b_{j})-h(\boldsymbol{\theta}^{\top}\boldsymbol{x})}\right|=\tilde{O}(N^{-1})$ (B.157)
with high probability. Finally, because $\|\boldsymbol{w}_{j}^{2T_{1}}-\boldsymbol{\theta}\|\leq\varepsilon$, we have
$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma\big{(}{\boldsymbol{w}_{j}^{2T_{1}}}^{\top}\boldsymbol{x}+b_{j}\big{)}-\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma(\boldsymbol{\theta}^{\top}\boldsymbol{x}+b_{j})}\right|=\tilde{O}\big{(}(\boldsymbol{w}_{j}^{2T_{1}}-\boldsymbol{\theta})^{\top}\boldsymbol{x}\big{)}=\tilde{O}(\varepsilon).$ (B.158)
Combining (B.157) and (B.158), we obtain the assertion. ∎

### v2

@@ B.6 Second Layer Training
$\!\!\!\!\!\!\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}$ $\lesssim(N^{-2}+\varepsilon^{2})+\frac{1}{T_{2}\lambda\delta_{0}}\mathopen{}\mathclose{{}\left(2N^{2}Q(R_{w},b,q^{\prime})+\mathbb{E}_{\boldsymbol{x}}[(f_{*}(\boldsymbol{x}))^{4}]}\right)+\frac{3\lambda}{2}\|\boldsymbol{a}^{*}\|^{2},$ (B.131)
with probability $1-4\delta_{0}$. Thus, since $\|\boldsymbol{a}^{*}\|^{2}=\tilde{O}(N)$, by setting $T_{2}=\tilde{\Theta}((N^{4}Q^{2}(R_{w},b,q^{\prime})+\mathbb{E}[f_{*}(\boldsymbol{x})^{4}]^{2})\varepsilon^{-4})$, and $N=\tilde{\Theta}(\varepsilon^{-1})$, we obtain that (B.119)$\lesssim\varepsilon^{2}$. ∎
Higher Generative Exponent Functions.
For general link functions, under Assumption 4 and the bounded fourth moment of the link function, we have the following counterpart of Lemma 20, which provides the formal statement of Proposition 10.
Lemma 21.
Suppose that $\mathbb{E}[\sigma_{*}(\boldsymbol{\theta}^{\top}x)^{4}]<\infty$ and Assumption 4 hold. Then, by setting $\lambda=\tilde{\Theta}\mathopen{}\mathclose{{}\left(\sqrt{\frac{N^{2}}{T_{2}\delta_{0}}}}\right)$ for some $\delta_{0}>0$, the ridge estimator $\hat{\boldsymbol{a}}$ satisfies
$\|f_{\hat{\boldsymbol{a}}}-f_{*}\|_{L^{2}(P_{x})}^{2}$ $\lesssim\varepsilon^{2}+\frac{1}{\sqrt{T_{2}\delta_{0}}}\mathopen{}\mathclose{{}\left(N^{2}C_{4}+\mathbb{E}_{\boldsymbol{x}}[(f_{*})^{4}]}\right)+\frac{1}{\sqrt{T_{2}\delta_{0}}}\|\boldsymbol{a}^{*}\|^{2},$ (B.132)
with probability $1-\delta_{0}$. By taking $T_{2}=\tilde{\Theta}((N^{4}+N^{2})\varepsilon^{-4})$, we have
$\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{a}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]\lesssim\varepsilon^{2}.$
Furthermore, applying Lemma 12 and 13 yields that, when $\sigma_{*}=\sum_{j=0}^{\infty}\alpha_{j}\mathsf{He}_{j}$ satisfies $\sum_{j=0}^{\infty}j^{2}j!\alpha_{j}^{2}$ and $\mathbb{E}[\sigma_{*}(\boldsymbol{\theta}^{\top}x)^{4}]$ are bounded, with a properly designed randomized activation in Lemma 12, by taking $N=\tilde{\Theta}(\varepsilon^{-7})$ and $T_{2}=\tilde{\Theta}(\varepsilon^{-32})$, Algorithm 1 yields
$\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{a}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]\lesssim\varepsilon^{2},$
with probability $1-o_{d}(1)$.
Proof.  The proof is identical to that of Lemma 20, with the difference being that we replace the bounded moment assumptions with $\mathbb{E}[\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})^{4}]<\infty$ or Assumption 4. ∎
Approximation Guarantee.
Note that for non-polynomial link function with generative exponent $p_{*}\geq 3$, the approximation error is already controlled in Assumption 4 based on [BBSS22, Lemma 4.4, 4.5] (using activation function with a ReLU component). If $\sigma_{*}$ is a degree-$q$ polynomial, we have the following approximation result using polynomial activation, which follows Lemmas 29 and 30 of [OSSW24a].
@@
$\mathopen{}\mathclose{{}\left|\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma_{j}\big{(}{\boldsymbol{w}_{j}^{2T_{1}}}^{\top}\boldsymbol{x}+b_{j}\big{)}-h(\boldsymbol{\theta}^{\top}\boldsymbol{x})}\right|=\tilde{O}(N^{-1}+\varepsilon)$ (B.133)
@@
Lemma 22 can be established from the following result in [OSSW24a].